OpenStudy (anonymous):

(x^2+11x+30) / (x + 6) Divide

OpenStudy (anonymous):

\[= (x+5)(x+6)/(x+6) = x+5\]

OpenStudy (anonymous):

cancel (x+6) from denominator and numerator ^_^

OpenStudy (anonymous):

A.) x^2 - 5x - 25 B.) x + 11/6 x + 5 C.) x + 5 D.) x + 6x + 5 So It's C?

OpenStudy (anonymous):

yes! ^_^

OpenStudy (anonymous):

Sweet thank you soooo much!!

OpenStudy (anonymous):

a pleasure ^_^